Rachel invested her savings in two investment funds. The amount she invested in Fund A was 4 times as much as the amount she invested in Fund B. Fund A returned a 7% profit and Fund B returned a 6% profit. How much did she invest in Fund B, if the total profit from the two funds together was $2720?
Answer by TimothyLamb(4379)

x = amount invested in A
y = amount invested in B
---
x = 4y
0.07x + 0.06y = 2720
---
put the system of linear equations into standard form
---
x - 4y = 0
0.07x + 0.06y = 2720
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = amount invested in A = 32000
y = amount invested in B = 8000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---